OpenStudy (anonymous):

what is the standard form of the equation of the line through (0,-3) with slope 2/5

OpenStudy (anonymous):

All you need to find the equation is a point and a slope, you have both of them. Now just plug it into this formula: \[y-y_1=m(x-x_1)\] And solve for y.

OpenStudy (anonymous):

\[(x_1, y_1) = (0, -3)\]

OpenStudy (anonymous):

the answer would be -2/5x+y+-3 is that correct

OpenStudy (anonymous):

Just make sure it's in the standard form, aka, solved for y. y=mx+b

OpenStudy (anonymous):

while m is the slope and b is the free term.

OpenStudy (anonymous):

y-(-3)=2/5(x-0) y+3=2/5x y=2/5x-3